Assume a consumer has $40 to spend and for both products the marginal utilities are shown in the following table: Quantity MU A MU B 1 35 80 2 20 40 3 12 18 Assume that each product sells for $10 per unit. a) How many units of each product will the consumer purchase? b) Assume the price of product B rises to $20 per unit. How will this consumer allocate her budget now? c) If the prices of both products rise to $20 per unit, what will be the budget allocation?
1
Expert's answer
2021-10-24T18:14:34-0400

Solution:

a.).  Utility maximization is where: MUA/PA = MUB/PB

The closest for both products is when they are producing 2 units each.

Product A = 2 units

Product B = 2 units

40 = 10A + 10B

40 = 10(2) + 10(2)

40 = 20 + 20

40 = 40

 

b.). The consumer will now purchase 1 unit of product B and 2 units of product A:

40 = 10A + 20B

40 = 10(2) + 20(1)

40 = 20 + 20

40 = 40

 

c.). If both the prices of the two products rises to $20 each, then the consumer will only afford 1 unit each of both products.

40 = 20A + 20B

40 = 20(1) + 20(1)

40 = 20 + 20

40 = 40